# Prediction of causal genes at GWAS loci with pleiotropic gene regulatory effects using sets of correlated instrumental variables

### v1

@@ Acknowledgements
T.M. acknowledges support from the Research Council of Norway (project number 312045). J.L.M.B. acknowledges support from the Swedish Research Council (2018-02529 and 2022-00734), the Swedish Heart Lung Foundation (2017-0265 and 2020-0207), the Leducq Foundation AteroGen (22CVD04) and PlaqOmics(18CVD02) consortia; the National Institute of Health-National Heart Lung Blood Institute (NIH/NHLBI, R01HL164577; R01HL148167; R01HL148239, R01HL166428, and R01HL168174), American Heart Association Transformational Project Award 19TPA34910021, and from the CMD AMP fNIH program. T.M. and J.L.M.B. acknowledge from the European Union’s Horizon Europe (European Innovation Council) programme (grant agreement number 101115381 to T.M. and J.L.M.B).
Supplementary Figures
@@
Figure S1: Causal effect estimation on simulated data of one dimensional systems. (A, C) Causal diagrams for the simulation of one exposure $X$ for an outcome $Y$, influenced by one instrument $E$ with variable instrument strengths $a$ (A), or influenced by $n\geq 2$ instruments $E_{1},\dots,E_{n}$ with instrument strengths $a_{i}$ (C). (B, D) Distribution of estimated causal effects for $X$ (true effect size $0.3$), showing distributions across 1,000 independently simulated datasets across a range of sample sizes under different simulation scenarios with varying instrument strengths. (G) Distribution of estimated causal effects for $X$ (true effect size $0.3$) assuming the false diagram in (F) for inference when the true diagram that generated the data is in (E).
@@
Figure S2: Causal effect estimation on simulated data for underdetermined and overdetermined systems. (A) Causal diagram for the simulation of two exposures $X_{1}$ and $X_{2}$ for an outcome $Y$, influenced by three instruments $E_{1},E_{2},E_{3}$ where the number of causal variants is smaller than the number of cis-eGenes and other variants in the locus are merely associated by LD. (B, C) Distribution of estimated causal effects for $X_{1}$ (B, true effect size $0.2$) and $X_{2}$ (C, true effect size $0.6$) in the graph from (A). (C) Causal diagram for the simulation of one exposure $X$ for an outcome $Y$, with $n\geq 2$ shared instruments $E_{1},\dots,E_{n}$ where the number of causal variants is greater than the number of cis-eGenes. (E) Distribution of estimated causal effects for $X$ (true effect size $0.3$) from the graph in (D). Panels B, C, and E show distributions across 1,000 independently simulated datasets across a range of sample sizes.
@@
Figure S3: Causal effect estimation and univariate MR methods comparison on simulated data of overdetermined systems. (A) Causal diagram for the simulation of exposure $MRPS6$ for an outcome $CAD$, with $10$ shared instruments. (B) Distribution of estimated causal effects for $MRPS6$ (true effect size $0.64$) comparing the estimators Least squares, GMM, TWMR, IVW (Inverse Variance Weighted), WM (Weighted Median), Maxlik (Maximum Likelihood), and MR-Egger, showing distributions across 1,000 independently simulated datasets across a range of sample sizes using normally distributed instruments with randomly generated covariances with real LD values from the MRPS6 locus.

### v2

@@ Acknowledgments
M.K., S.B., and T.M. acknowledge support from the Research Council of Norway (project number 312045). J.L.M.B. acknowledges support from the Swedish Research Council (2018-02529 and 2022-00734), the Swedish Heart Lung Foundation (2017-0265 and 2020-0207), the Leducq Foundation AteroGen (22CVD04) and PlaqOmics(18CVD02) consortia; the National Institute of Health-National Heart Lung Blood Institute (NIH/NHLBI, R01HL164577; R01HL148167; R01HL148239, R01HL166428, and R01HL168174), American Heart Association Transformational Project Award 19TPA34910021, and from the CMD AMP fNIH program. T.M. and J.L.M.B. acknowledge the European Union’s Horizon Europe (European Innovation Council) programme (grant agreement number 101115381). The funders had no role in study design, data collection and analysis, decision to publish, or preparation of the manuscript.
Supporting information
SI File.
Supplementary Information. Contains seven supplementary figures and nine pages of supplementary methods.
S1 Fig
@@
Causal effect estimation on simulated data of one dimensional systems. (A, C) Causal diagrams for the simulation of one exposure $X$ for an outcome $Y$, influenced by one instrument $E$ with variable instrument strengths $a$ (A), or influenced by $n\geq 2$ instruments $E_{1},\dots,E_{n}$ with instrument strengths $a_{i}$ (C). (B, D) Distribution of estimated causal effects for $X$ (true effect size $0.3$), showing distributions across 1,000 independently simulated datasets across a range of sample sizes under different simulation scenarios with varying instrument strengths. (G) Distribution of estimated causal effects for $X$ (true effect size $0.3$) assuming the false diagram in (F) for inference when the true diagram that generated the data is in (E).
S2 Fig
@@
Causal effect estimation on simulated data for underdetermined and overdetermined systems. (A) Causal diagram for the simulation of two exposures $X_{1}$ and $X_{2}$ for an outcome $Y$, influenced by three instruments $E_{1},E_{2},E_{3}$ where the number of causal variants is smaller than the number of cis-eGenes and other variants in the locus are merely associated by LD. (B, C) Distribution of estimated causal effects for $X_{1}$ (B, true effect size $0.2$) and $X_{2}$ (C, true effect size $0.6$) in the graph from (A). (C) Causal diagram for the simulation of one exposure $X$ for an outcome $Y$, with $n\geq 2$ shared instruments $E_{1},\dots,E_{n}$ where the number of causal variants is greater than the number of cis-eGenes. (E) Distribution of estimated causal effects for $X$ (true effect size $0.3$) from the graph in (D). Panels B, C, and E show distributions across 1,000 independently simulated datasets across a range of sample sizes.
S3 Fig
Conditional F-statistic estimation on simulated data of over-determined systems. (A) Distribution of conditional F-statistic using the Mendelian Randomization package (B) Distribution of conditional F-statistic using the MVMR package (C) Distribution of estimated causal effects for $ADAMTS7$ (true effect size $0.15$) for the estimator GMM, showing distributions across 2,000 independently simulated datasets across a range of sample sizes using discrete instruments with randomly generated covariances with real LD values from the ADAMTS7-CTSH locus on $15:79124475$ in the Aorta tissue. (D) Distribution of estimated causal effects for $ADAMTS7$ (true effect size $0.15$) for the estimator mvivw estimator using the Mendelian Randomization package which returns the estimates along with the conditional F-statistics.
S4 Fig
@@
Standard error estimation and causal effect estimation comparison on simulated data of overdetermined systems. (A) Distribution of estimated standard errors for $SLC22A3$, using the individual level data and exact form compared to using approximate form with summary level data. (B) Distribution of estimated causal effects for $SLC22A3$ (true effect size $0.15$) for the estimator GMM, showing distributions across 2,000 independently simulated datasets across a range of sample sizes using discrete instruments with randomly generated covariances with real LD values from the SLC22A3-LPA-PLG locus.
S5 Fig
Bias in univariate two-sample MR. Difference $(\hat{c}-c)$ between estimated and true causal effect of X (true effect size c = 0.8) on Y across 20,000 independently simulated datasets for a range of sample sizes comparing bias of two sample MR vs one sample MR, in a simple ratio estimate with one instrument E. (A) $\operatorname{Cov}(E,X)$ and $\operatorname{Cov}(E,Y)$ estimated from the same sample, Sample 1. (B) $\operatorname{Cov}(E,X)$ estimated from sample, Sample 1 and $\operatorname{Cov}(E,Y)$ estimated from sample, Sample 2. (C) $\operatorname{Cov}(E,X)$ estimated from sample, Sample 2 and $\operatorname{Cov}(E,Y)$ estimated from sample, Sample 1. (D) $\operatorname{Cov}(E,X)$ and $\operatorname{Cov}(E,Y)$ estimated from the same sample, Sample 2. For simplicity, sizes of Sample 1 and Sample 2 were kept the same.
S6 Fig
Standard error verification for STARNET. Tissue-wise causal estimates from the Least squares estimator (red) and MVIVW estimator from the Mendelian Randomization R package, with standard errors (blue).
S7 Fig
Estimated causal effects and their p-values. Tissue-wise causal estimates from the MVIVW estimator from the Mendelian Randomization R package, with their corresponding $-\log_{10}(p-value)$. One outlier (CDKN2B) with estimated effect size -1.0 and p-value $<10^{-80}$ not shown.